Problem 10-15 Return on Investment (ROI) and Residual Income [LO10-1, LO10-2]

Financial data for Joel de Paris, Inc., for last year, follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 140,000 $ 120,000
Accounts receivable 450,000 530,000
Inventory 320,000 380,000
Plant and equipment, net 680,000 620,000
Investment in Buisson, S.A. 250,000 280,000
Land (undeveloped) 180,000 170,000
Total assets $ 2,020,000 $ 2,100,000
Liabilities and Stockholders' Equity
Accounts payable $ 360,000 $ 310,000
Long-term debt 1,500,000 1,500,000
Stockholders' equity 160,000 290,000
Total liabilities and stockholders' equity $ 2,020,000 $ 2,100,000
Joel de Paris, Inc.
Income Statement
Sales $ 4,050,000
Operating expenses 3,645,000
Net operating income 405,000
Interest and taxes:
Interest expense $ 150,000
Tax expense 110,000 260,000
Net income $ 145,000


The company paid dividends of $15,000 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Turnover" to 1 decimal place.)

3. What was the company’s residual income last year?
Answer #1
Beginning balance Ending balance
cash 140000 120000
Accounts receivables 450000 530000
Inventory 320000 380000
Plant and equipment, net 680000 620000
Operating assets 1590000 1650000
Note : Investments in Buisson, S.A. and land (undeveloped) are not an operating assets as they do not contribute to the operating activities of the company.
1. Average operating assets = ( Beginning operating assets + Ending operating assets ) / 2 = ( 1590000 + 1650000 ) / 2 1620000
2. Margin = Net operating income / Sales = 405000 / 4050000 10%
Turnover = Sales / Average operating assets = 4050000 / 1620000 2.5
Return on investment (ROI) = Net operating income / Average operating assets = 405000 / 1620000 25%
3. Residual income = Net operating income - ( Minimum Required rate of return * Average operating assets ) = 405000 - ( 15% * 1620000 ) 162000